# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v3

@@ 1 Introduction
Gradient descent (GD) (Cauchy et al., 1847) and its stochastic variants are the preferred optimization methods in machine learning. The practical effectiveness of gradient methods heavily relies on the choice of the step-size (“learning rate”) parameter. Backtracking Armijo line-search (Armijo, 1966; Nocedal & Wright, 2006) (referred to as Armijo-LS) is a standard method to set the step-size for gradient descent. Given an initial step-size, the simplest form of Armijo-LS “searches” for the largest step-size that guarantees a sufficient decrease in the function value. When minimizing $L$-smooth functions using GD, Armijo-LS alleviates the need to know $L$, the global smoothness constant and enables setting the GD step-size in an adaptive manner. For both $L$-smooth convex and non-convex functions, GD with Armijo-LS (henceforth GD-LS) has been shown to match the favorable theoretical guarantees of GD with a constant $\nicefrac{{1}}{{L}}$ step-size (henceforth GD(1/L)). However, empirically, GD-LS typically results in faster convergence and is consequently, the default choice in practice.
One often-cited reason to explain the faster convergence of GD-LS is that it adapts to the “local” smoothness constant $L(\theta)$ near the point $\theta$, and results in an effective step-size of $\nicefrac{{1}}{{L(\theta)}}$. In some regions, $L(\theta)$ might be much smaller than the global smoothness $L$, thus allowing GD-LS to use much bigger step-sizes and consequently lead to faster convergence. Existing works that define notions of local smoothness (Scheinberg et al., 2014; Lu & Mei, 2023; Mishkin et al., 2024; Fox & Schmidt, , and references there in) can be used to formalize this intuition and show that GD-LS can result in constant factor improvements over GD(1/L), while having the same rate of convergence. In contrast, we consider a special class of objective functions and show that GD-LS (with a large initial step-size) can result in a provably faster rate of convergence compared to GD(1/L). In particular, we make the following contributions.
@@
Contribution 3. In Sec. 4, we instantiate Thm. 1 for non-uniform smooth, convex losses that include logistic regression and multi-class classification with the cross-entropy loss as examples. Specifically, in Cor. 1, we show that GD-LS (with a sufficiently large initial step-size) converges at an $O\left(\left(\nicefrac{{f^{*}}}{{\epsilon}}\right)\,\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate where $f^{*}:=\inf f(\theta)$. Hence, when $f^{*}$ is $\Theta(\epsilon)$, GD-LS converges at an $O\left(\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate, compared to the sublinear $O\left(\nicefrac{{1}}{{\epsilon}}\right)$ convergence of GD(1/L). We further instantiate this result for logistic regression on linearly separable data, and prove the linear convergence of GD-LS (Cor. 2), thus matching the rate for normalized GD (Axiotis & Sviridenko, 2023). In App. C, we show that GD with the Polyak step-size (Polyak, 1987) can inherit this fast convergence for logistic regression.
@@
Contribution 4. In Sec. 5, we instantiate Thm. 1 for non-convex functions satisfying non-uniform smoothness and gradient domination conditions that guarantee global optimality. Specifically, in Sec. 5.1, we analyze the convergence of GD-LS for the softmax policy gradient objective in reinforcement learning (Mei et al., 2020). In this setting, the linear convergence rate attained by GD-LS is provably better than the $\Omega(\nicefrac{{1}}{{\epsilon}})$ convergence of GD(1/L) and matches the rate of natural policy gradient (Kakade & Langford, 2002). In Sec. 5.2, we analyze the convergence of GD-LS for functions satisfying the PL condition (Polyak, 1987; Karimi et al., 2016) and instantiate the result for generalized linear models with the logistic link function. Our result demonstrates that GD-LS can converge faster than both constant step-size and normalized GD (Hazan et al., 2015).
Contribution 5. Finally, in Sec. 6, we investigate whether the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search proposed in Vaswani et al. (2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. The interpolation assumption is of practical interest (Zhang et al., 2017; Belkin et al., 2019), for example, it is satisfied for logistic regression with linearly separable data.
@@ 2 Problem Formulation
If $L_{1}=0$Assn. 2 recovers the standard uniform smoothness condition as a special case. Consequently, common smooth objectives such as linear regression or logistic regression satisfy the above condition. For example, if $X\in\mathbb{R}^{d\times n}$ is the feature matrix, and $y\in\mathbb{R}^{n}$ is the vector of measurements, then the linear regression objective, $f(\theta)=\frac{1}{2n}\,\left\|X\,\theta-y\right\|_{2}^{2}$ is $(\frac{1}{n}\lambda_{\max}[X^{T}X],0)$ non-uniform smooth where $\lambda_{\max}[A]$ is the maximum eigenvalue of the positive semi-definite matrix $A$.
In order to show the benefit of GD with Armijo line-search, we will focus on functions where $L_{1}\neq 0$, and the smoothness depends on the function value. We will require these functions to satisfy an additional assumption that relates the gradient norm to the function value. As we will see, such an assumption is true for losses with an exponential tail, even when using a 2 layer neural network (Taheri & Thrampoulidis, 2023; Wu et al., 2024).
@@ 3 GD with Armijo Line-search
where $c\in(0,1)$ is a tunable parameter. The complete pseudo-code is described in Alg. 1. The parameter $\beta$ controls the backtracking and is typically set to $0.9$, while the parameter $c$ is typically set to a small value such as $10^{-4}$ (Nocedal & Wright, 2006).
@@
When using GD-LS for minimizing $L$ uniformly-smooth functions (corresponding to $L_{0}\neq 0$ and $L_{1}=0$ in Assn. 2), $\eta_{t}$} is constrained to lie in the $\left[\min\left\{\eta_{\max},\frac{2\,(1-c)\,\beta}{L}\right\},\eta_{\max}\right]$ range (Nocedal & Wright, 2006). Note that this bound holds for all $L$ uniformly-smooth functions, does not require convexity, and guarantees that the backtracking line-search will terminate at a non-zero step-size. The parameter $c$ controls the “aggressiveness” of the algorithm; small $c$ values encourage a larger step-size. Hence, Armijo-LS can be seen as method to obtain a step-size proportional to $1/L$, without the knowledge of the global smoothness constant.
These bounds on the step-size can be used to derive the convergence rate for GD-LS. For example, for uniformly $L$-smooth and convex functions, the standard analysis shows that GD-LS converges to an optimum at an $O(1/T)$ rate (Nocedal & Wright, 2006). It thus matches the rate of GD with a constant step-size equal to $1/L$ (henceforth referred to as GD(1/L)). However, as alluded to in Sec. 1, Armijo-LS enables GD to adapt to the “local” smoothness $L(\theta_{t})$ (the smoothness around iterate $\theta_{t}$), and results in faster convergence both in theory (Scheinberg et al., 2014; Lu & Mei, 2023; Mishkin et al., 2024; Fox & Schmidt, ), and in practice. In contrast to these works that show constant factor improvements for GD-LS, we study non-uniform smooth functions satisfying Assn. 2 and 3, and aim to show that GD-LS can result in a faster rate of convergence.
We first show that, when minimizing non-uniform smooth functions satisfying Assn. 2 and 3, GD-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. Similar to the standard analysis (Nocedal & Wright, 2006), it is straightforward to relax this assumption. We prove the following lemma that lower-bounds the step-size returned by GD-LS.
@@
We now present a meta-theorem (proved in App. B) that characterizes this fast convergence for both convex and non-convex functions. It requires an additional condition which lower-bounds the gradient norm in terms of the function sub-optimality. In Sec. 4, we use convexity to satisfy this condition, whereas in Sec. 5, we use gradient domination to satisfy it. We also require that the step-size is not constrained by the initial choice, but rather by the properties of the function. This can be achieved by using a large $\eta_{\max}$, which is greater than $\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$ for all $t$, or by using a forward-tracking line-search to (approximately) return the largest step-size satisfying the Armijo condition (Fridovich-Keil & Recht, 2019). For conciseness, we express this requirement as $\eta_{\max}=\infty$. We note that it is straightforward to relax it and get an explicit dependence on $\eta_{\max}$, albeit at the cost of clarity in our theoretical results.
@@
We now split the proof into two cases: Case (1) when $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$ for all $t\in[T]$. In this case, $\lambda_{1}\,f(\theta_{t})\geq\lambda_{0}$. Using this relation with Eq. 6 and following the arguments in Axiotis & Sviridenko (2023, Theorem 5.2) allows us to complete the proof of Case (1). For Case (2), we follow the proof of the first case for iterations $t\in[\tau]$ for which $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and obtain a similar rate. This corresponds to Phase 1. with faster convergence. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation with Eq. 6 and following the standard proof for uniformly smooth functions completes the proof for Phase 2. which results in slower convergence. Putting the results for both phases together completes the proof of Case (2). ∎
@@ 4 GD-LS for Convex Losses
Figure 1: Comparing GD-LS with $c=1/2$ and $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on the ijcnn dataset (Chang & Lin, 2011). $f^{*}$ is small and GD-LS converges faster.
@@
In order to compare the result in Cor. 1 with existing works, consider the special case of logistic regression. In this case, GD-LS matches the rate for a variant of normalized gradient descent (NGD) in Axiotis & Sviridenko (2023, Theorem 5.2). However, unlike GD-LS, NGD requires the knowledge of $L_{1}$ making it relatively difficult to be implemented in practice. Furthermore, we note NGD is a specialized algorithm that is helpful to attain faster rates for certain problems (Mei et al., 2021; Hazan et al., 2015; Wilson et al., 2019), whereas GD-LS is universally used and can automatically exploit the problem structure. Moreover, Cor. 1 is more general and also holds for the exponential loss.
While GD(1/L) results in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 0Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
@@ 5 GD-LS for Non-convex Losses
Gradient domination or Łojasiewicz conditions are satisfied for matrix factorization (Ward & Kolda, 2023), policy gradient in reinforcement learning (Mei et al., 2020) and generalized linear models (Mei et al., 2021). These conditions have been exploited to prove global convergence guarantees for first-order methods (Karimi et al., 2016; Mei et al., 2021).

### v4

@@ 1 Introduction
Gradient descent (GD) (Cauchy and others, 1847) and its stochastic variants are the preferred optimization methods in machine learning. The practical effectiveness of gradient methods heavily relies on the choice of the step-size (“learning rate”) parameter. Backtracking Armijo line-search (Armijo, 1966; Nocedal and Wright, 2006) (referred to as Armijo-LS) is a standard method to set the step-size for gradient descent. Given an initial step-size, the simplest form of Armijo-LS “searches” for the largest step-size that guarantees a sufficient decrease in the function value. When minimizing $L$-smooth functions using GD, Armijo-LS alleviates the need to know $L$, the global smoothness constant and enables setting the GD step-size in an adaptive manner. For both $L$-smooth convex and non-convex functions, GD with Armijo-LS (henceforth GD-LS) has been shown to match the favorable theoretical guarantees of GD with a constant $\nicefrac{{1}}{{L}}$ step-size (henceforth GD(1/L)). However, empirically, GD-LS typically results in faster convergence and is consequently, the default choice in practice.
One often-cited reason to explain the faster convergence of GD-LS is that it adapts to the “local” smoothness constant $L(\theta)$ near the point $\theta$, and results in an effective step-size of $\nicefrac{{1}}{{L(\theta)}}$. In some regions, $L(\theta)$ might be much smaller than the global smoothness $L$, thus allowing GD-LS to use much bigger step-sizes and consequently lead to faster convergence. Existing works that define notions of local smoothness (Scheinberg et al., 2014; Lu and Mei, 2023; Mishkin et al., 2024; Fox and Schmidt, , and references there in) can be used to formalize this intuition and show that GD-LS can result in constant factor improvements over GD(1/L), while having the same rate of convergence. In contrast, we consider a special class of objective functions and show that GD-LS (with a large initial step-size) can result in a provably faster rate of convergence compared to GD(1/L). In particular, we make the following contributions.
@@
Contribution 3. In Sec. 4, we instantiate Thm. 1 for non-uniform smooth, convex losses that include logistic regression and multi-class classification with the cross-entropy loss as examples. Specifically, in Cor. 1, we show that GD-LS (with a sufficiently large initial step-size) converges at an $O\left(\left(\nicefrac{{f^{*}}}{{\epsilon}}\right)\,\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate where $f^{*}:=\inf f(\theta)$. Hence, when $f^{*}$ is $\Theta(\epsilon)$, GD-LS converges at an $O\left(\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ rate, compared to the sublinear $O\left(\nicefrac{{1}}{{\epsilon}}\right)$ convergence of GD(1/L). We further instantiate this result for logistic regression on linearly separable data, and prove the linear convergence of GD-LS (Cor. 2), thus matching the rate for normalized GD (Axiotis and Sviridenko, 2023). In App. C, we show that GD with the Polyak step-size (Polyak, 1987) can inherit this fast convergence for logistic regression.
@@
Contribution 4. In Sec. 5, we instantiate Thm. 1 for non-convex functions satisfying non-uniform smoothness and gradient domination conditions that guarantee global optimality. Specifically, in Sec. 5.1, we analyze the convergence of GD-LS for the softmax policy gradient objective in reinforcement learning (Mei et al., 2020). In this setting, the linear convergence rate attained by GD-LS is provably better than the $\Omega(\nicefrac{{1}}{{\epsilon}})$ convergence of GD(1/L) and matches the rate of natural policy gradient (Kakade and Langford, 2002). In Sec. 5.2, we analyze the convergence of GD-LS for functions satisfying the PL condition (Polyak, 1987; Karimi et al., 2016) and instantiate the result for generalized linear models with the logistic link function. Our result demonstrates that GD-LS can converge faster than both constant step-size and normalized GD (Hazan et al., 2015).
Contribution 5. Finally, in Sec. 6, we investigate whether the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search proposed in Vaswani et al. (2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt and Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. The interpolation assumption is of practical interest (Zhang et al., 2017; Belkin et al., 2019), for example, it is satisfied for logistic regression with linearly separable data.
@@ 2 Problem Formulation
If $L_{1}=0$, Assn. 2 recovers the standard uniform smoothness condition as a special case. Consequently, common smooth objectives such as linear regression or logistic regression satisfy the above condition. For example, if $X\in\mathbb{R}^{d\times n}$ is the feature matrix, and $y\in\mathbb{R}^{n}$ is the vector of measurements, then the linear regression objective, $f(\theta)=\frac{1}{2n}\,\left\|X\,\theta-y\right\|_{2}^{2}$ is $(\frac{1}{n}\lambda_{\max}[X^{T}X],0)$ non-uniform smooth where $\lambda_{\max}[A]$ is the maximum eigenvalue of the positive semi-definite matrix $A$.
In order to show the benefit of GD with Armijo line-search, we will focus on functions where $L_{1}\neq 0$, and the smoothness depends on the function value. We will require these functions to satisfy an additional assumption that relates the gradient norm to the function value. As we will see, such an assumption is true for losses with an exponential tail, even when using a 2 layer neural network (Taheri and Thrampoulidis, 2023; Wu et al., 2024).
@@ 3 GD with Armijo Line-search
where $c\in(0,1)$ is a tunable parameter. The complete pseudo-code is described in Alg. 1. The parameter $\beta$ controls the backtracking and is typically set to $0.9$, while the parameter $c$ is typically set to a small value such as $10^{-4}$ (Nocedal and Wright, 2006).
@@
When using GD-LS for minimizing $L$ uniformly-smooth functions (corresponding to $L_{0}\neq 0$ and $L_{1}=0$ in Assn. 2), $\eta_{t}$} is constrained to lie in the $\left[\min\left\{\eta_{\max},\frac{2\,(1-c)\,\beta}{L}\right\},\eta_{\max}\right]$ range (Nocedal and Wright, 2006). Note that this bound holds for all $L$ uniformly-smooth functions, does not require convexity, and guarantees that the backtracking line-search will terminate at a non-zero step-size. The parameter $c$ controls the “aggressiveness” of the algorithm; small $c$ values encourage a larger step-size. Hence, Armijo-LS can be seen as method to obtain a step-size proportional to $1/L$, without the knowledge of the global smoothness constant.
These bounds on the step-size can be used to derive the convergence rate for GD-LS. For example, for uniformly $L$-smooth and convex functions, the standard analysis shows that GD-LS converges to an optimum at an $O(1/T)$ rate (Nocedal and Wright, 2006). It thus matches the rate of GD with a constant step-size equal to $1/L$ (henceforth referred to as GD(1/L)). However, as alluded to in Sec. 1, Armijo-LS enables GD to adapt to the “local” smoothness $L(\theta_{t})$ (the smoothness around iterate $\theta_{t}$), and results in faster convergence both in theory (Scheinberg et al., 2014; Lu and Mei, 2023; Mishkin et al., 2024; Fox and Schmidt, ), and in practice. In contrast to these works that show constant factor improvements for GD-LS, we study non-uniform smooth functions satisfying Assn. 2 and 3, and aim to show that GD-LS can result in a faster rate of convergence.
We first show that, when minimizing non-uniform smooth functions satisfying Assn. 2 and 3, GD-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. Similar to the standard analysis (Nocedal and Wright, 2006), it is straightforward to relax this assumption. We prove the following lemma that lower-bounds the step-size returned by GD-LS.
@@
We now present a meta-theorem (proved in App. B) that characterizes this fast convergence for both convex and non-convex functions. It requires an additional condition which lower-bounds the gradient norm in terms of the function sub-optimality. In Sec. 4, we use convexity to satisfy this condition, whereas in Sec. 5, we use gradient domination to satisfy it. We also require that the step-size is not constrained by the initial choice, but rather by the properties of the function. This can be achieved by using a large $\eta_{\max}$, which is greater than $\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$ for all $t$, or by using a forward-tracking line-search to (approximately) return the largest step-size satisfying the Armijo condition (Fridovich-Keil and Recht, 2019). For conciseness, we express this requirement as $\eta_{\max}=\infty$. We note that it is straightforward to relax it and get an explicit dependence on $\eta_{\max}$, albeit at the cost of clarity in our theoretical results.
@@
We now split the proof into two cases: Case (1) when $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$ for all $t\in[T]$. In this case, $\lambda_{1}\,f(\theta_{t})\geq\lambda_{0}$. Using this relation with Eq. 6 and following the arguments in Axiotis and Sviridenko (2023, Theorem 5.2) allows us to complete the proof of Case (1). For Case (2), we follow the proof of the first case for iterations $t\in[\tau]$ for which $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and obtain a similar rate. This corresponds to Phase 1. with faster convergence. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation with Eq. 6 and following the standard proof for uniformly smooth functions completes the proof for Phase 2. which results in slower convergence. Putting the results for both phases together completes the proof of Case (2). ∎
@@ 4 GD-LS for Convex Losses
Figure 1: Comparing GD-LS with $c=1/2$ and $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on the ijcnn dataset (Chang and Lin, 2011). $f^{*}$ is small and GD-LS converges faster.
@@
In order to compare the result in Cor. 1 with existing works, consider the special case of logistic regression. In this case, GD-LS matches the rate for a variant of normalized gradient descent (NGD) in Axiotis and Sviridenko (2023, Theorem 5.2). However, unlike GD-LS, NGD requires the knowledge of $L_{1}$ making it relatively difficult to be implemented in practice. Furthermore, we note NGD is a specialized algorithm that is helpful to attain faster rates for certain problems (Mei et al., 2021; Hazan et al., 2015; Wilson et al., 2019), whereas GD-LS is universally used and can automatically exploit the problem structure. Moreover, Cor. 1 is more general and also holds for the exponential loss.
While GD(1/L) results in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov and others, 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 0. Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
@@ 5 GD-LS for Non-convex Losses
Gradient domination or Łojasiewicz conditions are satisfied for matrix factorization (Ward and Kolda, 2023), policy gradient in reinforcement learning (Mei et al., 2020) and generalized linear models (Mei et al., 2021). These conditions have been exploited to prove global convergence guarantees for first-order methods (Karimi et al., 2016; Mei et al., 2021).